# Identifying Constitutive Parameters for Complex Hyperelastic Materials using Physics-Informed Neural Networks

### v3

@@ 1 Introduction
The emerging demands to characterize biological materials and design advanced bio-inspired metamaterials necessitate precisely identifying their material properties, such as Young’s modulus and shear modulus. Traditionally, the modulus is identified through mechanical tests Davis (2004), such as uniaxial tensile experiments performed on specially prepared samples that comply with testing protocols. This approach involves collecting stress-strain or pressure-volume relationships Jin et al. (2021) and presuming a phenomenological constitutive model for the material, such as the Neo-Hookean (NH) model Treloar (1948). Subsequently, the material modulus can be identified by fitting the experimental data with the presumed model. As the complexity of the constitutive model increases (e.g., Fung-type model Fung et al. (1979), Holzapfel-Gasser-Ogden (HGO) model Holzapfel et al. (2000)), a variety of testing methodologies, including shear and biaxial testing, become necessary to accurately identify the constitutive parameters Sacks (2000); Sacks and Sun (2003). Nevertheless, this traditional framework requires a strict sample preparation process. For example, samples should have a carefully designed dog-bone shape to allow parameter identification under uniaxial testing. However, shaping certain biological tissues like cruciate ligaments into such specific geometry proves difficult Luetkemeyer et al. (2021). Furthermore, the advent of novel materials like mechanical metamaterials Jin and Espinosa (2024); Bertoldi et al. (2017), which contain intentionally designed defects, makes conventional methodologies, like finite element model update methods or virtual field methods Avril et al. (2008), insufficient for identifying their homogenized material parameters. Consequently, identifying the material parameters for samples with intricate geometries comprising irregular boundaries and internal defects continues to have challenges.
@@
Recent advances in deep learning have seen a surge in the application of end-to-end material properties identifications Gu et al. (2018); Jin et al. (2022); Ni and Gao (2021); Bock et al. (2019); Yang et al. (2021); Jin (2022); Guo et al. (2021); Liu et al. (2020); Yang et al. (2020); Jin et al. (2023a). In most cases, these methods require a sufficient training dataset, sourced primarily from simulations. However, this approach often requires significant computations, particularly when working with high-dimensional parameter spaces Jin et al. (2023b). While traditional data-driven methods embed fundamental physics, such as governing equations, into simulations, there is an increasing demand for more effective methodologies. The focus has shifted towards methods that directly encode these physical laws into deep learning algorithms, enhancing the efficiency of material modulus identification. To this end, Physics-Informed Neural Networks (PINNs) Raissi et al. (2019); Karniadakis et al. (2021) have been developed to explicitly encode the underlying physical laws, i.e., partial differential equations (PDEs), into the neural networks (NNs). This integration of prior physical law knowledge into a deep learning framework can notably minimize the required size of the training dataset without compromising prediction accuracy. Since their inception, PINNs have demonstrated substantial utility across a broad spectrum of engineering disciplines, including solid mechanics Haghighat et al. (2021); Zhang et al. (2020, 2022); Li et al. (2021); Bastek and Kochmann (2023); Chen and Gu (2023); Henkes et al. (2022); Niu et al. (2023), fluid mechanics Cai et al. (2021); Raissi et al. (2020); Cai et al. (2021); Jin et al. (2021), and biomaterials Kamali et al. (2023); Yin et al. (2021). As an example of the material modulus identifications, Zhang et al. Zhang et al. (2020, 2022) demonstrated that PINNs could efficiently identify the distribution of inhomogeneous material and geometry under the plane strain condition, bypassing the need for computation-intensive data-driven processes. These PINNs frameworks, specifically tailored for continuum solid mechanics, hold promising potential to address the identified challenges in modulus identification of solids with intricate constitutive models. As such, they represent an exciting frontier in material properties identification. However, it is important to note that the current PINNs frameworks still have their limitations, with their capacity largely restricted to identifying parameters from simple constitutive laws such as linear elasticity Kamali et al. (2023); Anton and Wessels (2022); Chen and Gu (2023) and NH solids Zhang et al. (2020). Recent endeavors by Hamel et al. Hamel et al. (2023) seek to enhance the original PINNs framework to calibrate constitutive parameters for complex hyperelastic models using full-field measurements. However, the adoption of a computational grid to integrate the PDE’s weak form poses challenges when combined with experimental full-field measurement such as Digital Image Correlation (DIC). Therefore, a robust and applicable PINN-based algorithm for identifying constitutive parameters of complex hyperelastic solids is still lacking.
@@
Herein, we introduce a PINN-based deep learning framework, which can accurately identify material parameters for complex constitutive behaviors beyond NH from multi-modal synthetic experimental data. Compared to the original framework Zhang et al. (2022), we integrated the complete deformation and loading history into the PINNs under the plane stress condition, which is more practical to the experimental setup. To accommodate inhomogeneous deformation, particularly when substantial internal defects are present, we coupled the full-field displacements measured from DIC Chu et al. (1985) as additional boundary conditions (BCs) for training. These full-field displacement measurements provide valuable local deformation data induced by internal inhomogeneities. Therefore, by combining the full-field displacement field into the PINNs framework, we can effectively predict material characteristics even in scenarios where conventional methods prove insufficient due to the complexity of biological samples. In Section 2, we will briefly review the fundamentals of PINNs for continuum solid mechanics and present our detailed algorithm. In Section 3, we will first verify our framework by solving a forward problem and compare the results with those obtained from finite element methods (FEM). Subsequently, we will employ the framework to identify the constitutive parameters for Arruda-Boyce (AB) constitutive laws using benchmark problems: rectangular samples with circular inhomogeneities inside. We will also discuss the quality and noise effects of DIC data on the overall accuracy and efficiency of prediction. We anticipate that our framework will provide a robust modulus identification approach for soft materials with complex constitutive behaviors and geometries using direct experimental measurements.
@@ 2 PINNs framework for parameter identification
The methodology for data collection and processing for our algorithm is illustrated in Fig.1(a). A sample with internal geometric anomalies and unidentified material parameters is subjected to uniaxial tension under the plane stress condition. The sample’s top surface is speckled and subsequently captured with a digital camera. DIC analysis Chu et al. (1985) can then be employed to obtain the full-field displacements. Simultaneously, a load cell records the loading under deformation. To the end, a multi-modal experimental dataset can be collected, containing both the full-field displacement field and loading history. This dataset subsequently serves as additional BCs for the PINNs training, aiding in identifying constitutive parameters. In Section 2.1, we first introduce the PINNs framework, tailored for continuum solids with hyperelastic constitutive behaviors under large deformation. Then, we encompass the setup for a benchmark problem in Section 2.2, utilized to validate our framework via both forward surrogate modeling and inverse parameter identification. In this paper, we used FEM to generate the synthetic multi-modal experimental data to validate our framework.
@@ 2.1 PINNs for continuum solid mechanics
Here, we briefly review PINNs for continuum solid mechanics and introduce our modifications for complex hyperelastic solids beyond simple NH materials under the quasi-static plane stress deformation. The readers can refer to recent research papers Raissi et al. (2019, 2020); Zhang et al. (2022) and reviews Jin et al. (2023b); Cai et al. (2021) for detailed descriptions of PINNs. The architecture of PINNs is depicted in Fig.1(b). For convenience, we denote directions 1, 2, and 3 as the loading direction, transverse direction, and out-of-plane direction, respectively. We define the unknown constitutive parameters as $\bm{\theta}_{\text{const}}$. There are in total four key steps to identify $\bm{\theta}_{\text{const}}$ in the PINNs framework.
@@
Second, we integrate continuum mechanics into the PINNs by deriving relevant quantities from the NN outputs. During the training process, the partial derivatives are processed by automatic differentiation Baydin et al. (2018). Therefore, the deformation gradient $\tilde{\textbf{F}}(\textbf{X};t)$ and the first Piola-Kirchhoff stress $\tilde{\textbf{P}}(\textbf{X};t)$ of a generalized hyperelastic solid can be calculated as,
@@ 2.2 Setup of the prototype problem
where $N$ is the number of chains in polymer network, $n$ is the number of chain segments, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. Here, $\lambda_{\text{chain}}=\sqrt{\frac{I_{1}}{3}}$ with $I_{1}$ the first invariant of left Cauchy-Green tensor. $\beta=\mathcal{L}^{-1}(\frac{\lambda_{\text{chain}}}{\sqrt{n}})$ with $\mathcal{L}^{-1}(\bullet)$ the inverse Langevin function. If we consider the first five terms of the inverse Langevin function, the strain energy density can be approximated as Systèmes (2021),
@@ 2.3 Training of PINNs
We implemented and customized the PINNs framework on the open-source Python library, Deepxde 1.7.0, developed by Lu et al. Lu et al. (2021). The PINNs architecture has a total of 3 hidden layers with 30 neurons in each layer. In addition, we set Tanh as the activation function and Glorot uniform as the weight initializer. Through this paper, we set Adam Kingma and Ba (2014) as the optimization algorithm with a learning rate of 0.001. Other higher-order optimization algorithms like L-BFGS can be also employed Liu and Nocedal (1989) to further improve training efficiency. The PINNs were trained on a laptop with a 14-core i7-12700H CPU and a RTX3060 GPU. Leveraging GPU acceleration, we achieved a training speed of approximately 50 epochs per second.
@@ 2.4 Finite Element Analysis
We performed the FEM using Abaqus 2021 Static package Systèmes (2021) to verify the forward problems and also generate the synthetic DIC and loading history dataset for the inverse problem. The medium was modeled using the incompressible AB constitutive law. The left boundary was clamped, while the right boundary had a displacement BC to deform the sample. The simulation domain was meshed with four-node plane stress elements (CPS4R). The linear mesh density is approximately 111 per unit length and mesh refinement near the internal defects was performed. Mesh dependence of the loading curve and the displacement field was examined, and all solutions reported are final converged solutions. The loading history was obtained by integrating the total force on the right boundary, and synthetic DIC data was extracted from the displacement field at the FEM nodes. We applied a Gaussian random noise with the desired level for the cases considering the DIC noise effects.
@@ 3 Results
where $\Lambda$ is the stretch ratio. The resulting $(x_{1},x_{2})$ inherently satisfies the displacement BCs Lu et al. (2021); Bai et al. (2023). This transformation can reduce the total number of loss functions during training and also ensure that the trivial solution, $f,g=0$, is close to the theoretical solution, leading to the reduction of the error in the power term $\frac{I_{1}^{n}}{\lambda_{m}^{2n}}(n=1,2,3,...)$. Furthermore, we consider the following unitless material and geometry parameters, $\lambda_{m}=3$, $\mu=1$, $a=1$, $b=0.5$, and the maximum stretch $\Lambda_{max}=2.5$. In biomechanics literature, the strain hardening parameter $\lambda_{m}$ of the AB model is measured as $1.6$ for brain white tissues Ashrafi and Shariyat (2017), and $3.32\pm 3.56$ for fibroadenoma O’Hagan and Samani (2009). The selected value $\lambda_{m}=3$ is within the typical range for soft tissues. Through this paper, we normalized the loss functions by dividing them by the total loss at the initial step.
@@ 3.1 Forward Problem
The performance for material identification depends on the PINNs’ capability to accurately predict the deformation field of the sample. To this end, in this forward problem, we investigate a rectangular domain featuring a centrally located hole with a radius of $r=0.15$, as illustrated in Fig.2(a). The parameters of PINNs, $\bm{\theta}_{NN}$, were optimized by minimizing the loss functions of incompressibility $\mathcal{L}^{\text{inc}}$, equilibrium equations $\mathcal{L}^{\text{PDE}}$, displacement BCs $\mathcal{L}^{\text{BC1}}$ at $X_{1}=\pm 0.5$, and traction-free BCs $\mathcal{L}^{\text{BC2}}$ at $r=0.15$ and $X_{2}=\pm 0.5$. The associated weights for these loss functions were tuned to yield the best performance. Our training dataset comprised 20,000 randomly sampled points within the time-spatial domain, supplemented by 50 uniformly sampled points within the hole region and 200 uniformly sampled points along the free boundary, spanning across 21 evenly spaced time steps. We trained the PINNs for 1,000k epochs.
@@
Figure 2(b) shows the losses rapidly reach $10^{-3}$ after 200k training epochs and converge to $10^{-3}-10^{-4}$ after 1, [^0] K training epochs, which indicates our PINNs can optimize the NN parameters with incorporated physics efficiently. The deformation outlines under five different stretch ratios, as depicted in Fig.2(c), showcase a strong agreement between predictions from the PINNs and the FEM results. Furthermore, the normalized loading forces predicted from PINNs also agree well with the FEM (Fig.2(d)), demonstrating the strain-stiffening effect in most elastomers with chemical cross-links, which is attributed to the parameter $\lambda_{m}$ in the AB model. To demonstrate the prediction of stress field from PINNs, we plotted the Cauchy stress contour in the loading direction, $\sigma_{11}$, in Fig.2(e). We found that the Cauchy stress contours agree well between the PINNs and FEM far away from the hole. While the relative error distanced from the hole remains small ($10^{-2}-10^{-3}$), large errors emerge in the proximity of the hole, with an average value of 0.6. This discrepancy is primarily attributed to a stress concentration around the hole, contributing to a decaying field characterized by a substantial displacement gradient Bower (2009); Anand and Govindjee (2020), especially when the sample is under large deformation. To enhance the accuracy within this localized region, it becomes imperative to employ a multiscale approach capable of accurately representing the local decaying field. For example, combining deep neural operator (DeepONet) Lu et al. (2021) for local surrogate modeling with global FEM solver has recently shown a promising solution to increase the local simulation accuracy Yin et al. (2022). However, it is worth noting that our domain integration method effectively captures the long-range effect of the hole on the loading curve (Fig.2(d)). Therefore, we believe our PINNs framework demonstrates noteworthy precision in assessing both the displacement field and the loading history, and the localized stress concentration will have negligible effects on parameter identification in the inverse problem, to be shown in the following subsections.
3.2 Identify constitutive parameters for the phototype problem

### v4

@@ 1 Introduction
The emerging demands to characterize biological materials and design advanced bio-inspired metamaterials necessitate precisely identifying their material properties, such as Young’s modulus and shear modulus. Traditionally, the modulus is identified through mechanical tests [1], such as uniaxial tensile experiments performed on specially prepared samples that comply with testing protocols. This approach involves collecting stress-strain or pressure-volume relationships [2] and presuming a phenomenological constitutive model for the material, such as the Neo-Hookean (NH) model [3]. Subsequently, the material modulus can be identified by fitting the experimental data with the presumed model. As the complexity of the constitutive model increases (e.g., Fung-type model [4], Holzapfel-Gasser-Ogden (HGO) model [5]), a variety of testing methodologies, including shear and biaxial testing, become necessary to accurately identify the constitutive parameters [6, 7]. Nevertheless, this traditional framework requires a strict sample preparation process. For example, samples should have a carefully designed dog-bone shape to allow parameter identification under uniaxial testing. However, shaping certain biological tissues like cruciate ligaments into such specific geometry proves difficult [8]. Furthermore, the advent of novel materials like mechanical metamaterials [9, 10], which contain intentionally designed defects, makes conventional methodologies, like finite element model update methods or virtual field methods [11], insufficient for identifying their homogenized material parameters. This inadequacy primarily arises from the inherent difficulties associated with measuring internal displacement fields in such complex structures. Consequently, identifying the material parameters for samples with intricate geometries comprising irregular boundaries and internal defects continues to have challenges.
@@
Recent advances in deep learning have seen a surge in the application of end-to-end material properties identifications [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In most cases, these methods require a sufficient training dataset, sourced primarily from simulations. However, this approach often requires significant computations, particularly when working with high-dimensional parameter spaces [22]. While traditional data-driven methods embed fundamental physics, such as governing equations, into simulations, there is an increasing demand for more effective methodologies. The focus has shifted towards methods that directly encode these physical laws into deep learning algorithms, enhancing the efficiency of material modulus identification. To this end, Physics-Informed Neural Networks (PINNs) [23, 24] have been developed to explicitly encode the underlying physical laws, i.e., partial differential equations (PDEs), into the neural networks (NNs). This integration of prior physical law knowledge into a deep learning framework can notably minimize the required size of the training dataset without compromising prediction accuracy. Since their inception, PINNs have demonstrated substantial utility across a broad spectrum of engineering disciplines, including solid mechanics [25, 26, 27, 28, 29, 30, 31, 32], fluid mechanics [33, 34, 35, 36], and biomaterials [37, 38]. As an example of the material modulus identifications, Zhang et al. [26, 27] demonstrated that PINNs could efficiently identify the distribution of inhomogeneous material and geometry under the plane strain condition, bypassing the need for computation-intensive data-driven processes. These PINNs frameworks, specifically tailored for continuum solid mechanics, hold promising potential to address the identified challenges in modulus identification of solids with intricate constitutive models. As such, they represent an exciting frontier in material properties identification. However, it is important to note that the current PINNs frameworks still have their limitations, with their capacity largely restricted to identifying parameters from simple constitutive laws such as linear elasticity [37, 39, 30] and NH solids [26]. Recent endeavors by Hamel et al. [40] seek to enhance the original PINNs framework to calibrate constitutive parameters for complex hyperelastic models using full-field measurements. However, the adoption of a computational grid to integrate the PDE’s weak form poses challenges when combined with experimental full-field measurement such as Digital Image Correlation (DIC). Therefore, a robust and applicable PINN-based algorithm for identifying constitutive parameters of complex hyperelastic solids is still lacking.
@@
Herein, we introduce a PINN-based deep learning framework, which can accurately identify material parameters for complex constitutive behaviors beyond NH from multi-modal synthetic experimental data. Compared to the original framework [27], we integrated the complete deformation and loading history into the PINNs under the plane stress condition, which is more practical to the experimental setup. To accommodate inhomogeneous deformation, particularly when substantial internal defects are present, we coupled the full-field displacements measured from DIC [41] as additional boundary conditions (BCs) for training. These full-field displacement measurements provide valuable local deformation data induced by internal inhomogeneities. Therefore, by combining the full-field displacement field into the PINNs framework, we can effectively predict material characteristics even in scenarios where conventional methods prove insufficient due to the complexity of biological samples. In Section 2, we will briefly review the fundamentals of PINNs for continuum solid mechanics and present our detailed algorithm. In Section 3, we will first verify our framework by solving a forward problem and compare the results with those obtained from finite element methods (FEM). Subsequently, we will employ the framework to identify the constitutive parameters for Arruda-Boyce (AB) constitutive laws using prototype problems: rectangular samples with circular inhomogeneities inside. We will also discuss the quality and noise effects of DIC data on the overall accuracy and efficiency of prediction. We anticipate that our framework will provide a robust modulus identification approach for soft materials with complex constitutive behaviors and geometries using direct experimental measurements.
@@ 2 PINNs framework for parameter identification
The methodology for data collection and processing for our algorithm is illustrated in Fig.1(a). A sample with internal geometric anomalies and unidentified material parameters is subjected to uniaxial tension under the plane stress condition. The sample’s top surface is speckled and subsequently captured with a digital camera. DIC analysis [41] can then be employed to obtain the full-field displacements. Simultaneously, a load cell records the loading under deformation. To the end, a multi-modal experimental dataset can be collected, containing both the full-field displacement field and loading history. This dataset subsequently serves as additional BCs for the PINNs training, aiding in identifying constitutive parameters. In Section 2.1, we first introduce the PINNs framework, tailored for continuum solids with hyperelastic constitutive behaviors under large deformation. Then, we encompass the setup for a prototype problem in Section 2.2, utilized to validate our framework via both forward surrogate modeling and inverse parameter identification. In this paper, we used FEM to generate the synthetic multi-modal experimental data to validate our framework.
@@ 2.1 PINNs for continuum solid mechanics
Here, we briefly review PINNs for continuum solid mechanics and introduce our modifications for complex hyperelastic solids beyond simple NH materials under the quasi-static plane stress deformation. The readers can refer to recent research papers [23, 34, 27] and reviews [22, 33] for detailed descriptions of PINNs. The architecture of PINNs is depicted in Fig.1(b). For convenience, we denote directions 1, 2, and 3 as the loading direction, transverse direction, and out-of-plane direction, respectively. We define the unknown constitutive parameters as $\bm{\theta}_{\text{const}}$. There are in total four key steps to identify $\bm{\theta}_{\text{const}}$ in the PINNs framework.
@@
Second, we integrate continuum mechanics into the PINNs by deriving relevant quantities from the NN outputs. During the training process, the partial derivatives are processed by automatic differentiation [42]. Therefore, the deformation gradient $\tilde{\textbf{F}}(\textbf{X};t)$ and the first Piola-Kirchhoff stress $\tilde{\textbf{P}}(\textbf{X};t)$ of a generalized hyperelastic solid can be calculated as,
@@ 2.2 Setup of the prototype problem
where $N$ is the number of chains in polymer network, $n$ is the number of chain segments, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. Here, $\lambda_{\text{chain}}=\sqrt{\frac{I_{1}}{3}}$ with $I_{1}$ the first invariant of left Cauchy-Green tensor. $\beta=\mathcal{L}^{-1}(\frac{\lambda_{\text{chain}}}{\sqrt{n}})$ with $\mathcal{L}^{-1}(\bullet)$ the inverse Langevin function. If we consider the first five terms of the inverse Langevin function, the strain energy density can be approximated as [44],
@@ 2.3 Training of PINNs
We implemented and customized the PINNs framework on the open-source Python library, Deepxde 1.7.0, developed by Lu et al. [45]. The PINNs architecture has a total of 3 hidden layers with 30 neurons in each layer. In addition, we set Tanh as the activation function and Glorot uniform as the weight initializer. Through this paper, we set Adam [46] as the optimization algorithm with a learning rate of 0.001. Other higher-order optimization algorithms like L-BFGS can be also employed [47] to further improve training efficiency. The PINNs were trained on a laptop with a 14-core i7-12700H CPU and a RTX3060 GPU. Leveraging GPU acceleration, we achieved a training speed of approximately 50 epochs per second.
@@ 2.4 Finite Element Analysis
We performed the FEM using Abaqus 2021 Static package [44] to verify the forward problems and also generate the synthetic DIC and loading history dataset for the inverse problem. The medium was modeled using the incompressible AB constitutive law. The left boundary was clamped, while the right boundary had a displacement BC to deform the sample. The simulation domain was meshed with four-node plane stress elements (CPS4R). The linear mesh density is approximately 111 per unit length and mesh refinement near the internal defects was performed. Mesh dependence of the loading curve and the displacement field was examined, and all solutions reported are final converged solutions. The loading history was obtained by integrating the total force on the right boundary, and synthetic DIC data was extracted from the displacement field at the FEM nodes. We applied a Gaussian random noise with the desired level for the cases considering the DIC noise effects.
@@ 3 Results
where $\Lambda$ is the stretch ratio, $f$ and $g$ represent the variables that serve as intermediate outputs from the NNs. The resulting $(x_{1},x_{2})$ inherently satisfies the displacement BCs [45]. This transformation can reduce the total number of loss functions during training and also ensure that the trivial solution, $f,g=0$, is close to the theoretical solution, leading to the reduction of the error in the power term $\frac{I_{1}^{n}}{\lambda_{m}^{2n}}(n=1,2,3,...)$. Furthermore, we consider the following unitless material and geometry parameters, $\lambda_{m}=3$, $\mu=1$, $a=1$, $b=0.5$, and the maximum stretch $\Lambda_{max}=2.5$. In biomechanics literature, the strain hardening parameter $\lambda_{m}$ of the AB model is measured as $1.6$ for brain white tissues [48], and $3.32\pm 3.56$ for fibroadenoma [49]. The selected value $\lambda_{m}=3$ is within the typical range for soft tissues. Through this paper, we normalized the loss functions by dividing them by the total loss at the initial step.
@@ 3.1 Forward Problem
The performance for material identification depends on the PINNs’ capability to accurately predict the deformation field of the sample. To this end, in this forward problem, we investigate a rectangular domain featuring a centrally located hole with a radius of $r=0.15$, as illustrated in Fig.2(a). The parameters of PINNs, $\bm{\theta}_{NN}$, were optimized by minimizing the loss functions of incompressibility $\mathcal{L}^{\text{inc}}$, equilibrium equations $\mathcal{L}^{\text{PDE}}$, displacement BCs $\mathcal{L}^{\text{BC1}}$ at $X_{1}=\pm 0.5$, and traction-free BCs $\mathcal{L}^{\text{BC2}}$ at $r=0.15$ and $X_{2}=\pm 0.5$. The associated weights for these loss functions were tuned to yield the best performance. Our training dataset comprised 20,000 randomly sampled points within the time-spatial domain, supplemented by 50 uniformly sampled points along the hole boundary and 200 uniformly sampled points along the free boundary, spanning across 21 evenly spaced time steps. We trained the PINNs for 1,000k epochs.
@@
Fig.2(b) shows the losses rapidly reach $10^{-3}$ after 200k training epochs and converge to $10^{-3}-10^{-4}$ after 1, [^0] K training epochs, which indicates our PINNs can optimize the NN parameters with incorporated physics efficiently. The deformation outlines under five different stretch ratios, as depicted in Fig.2(c), showcase a strong agreement between predictions from the PINNs and the FEM results. Furthermore, the normalized loading forces predicted from PINNs also agree well with the FEM (Fig.2(d)), demonstrating the strain-stiffening effect in most elastomers with chemical cross-links, which is attributed to the parameter $\lambda_{m}$ in the AB model. To demonstrate the prediction of stress field from PINNs, we plotted the Cauchy stress contour in the loading direction, $\sigma_{11}$, in Fig.2(e). We found that the Cauchy stress contours agree well between the PINNs and FEM far away from the hole. While the relative error distanced from the hole remains small ($10^{-2}-10^{-3}$), large errors emerge in the proximity of the hole, with an average value of 0.6. This discrepancy is primarily attributed to a stress concentration around the hole, contributing to a decaying field characterized by a substantial displacement gradient [50, 51], especially when the sample is under large deformation. To enhance the accuracy within this localized region, it becomes imperative to employ a multiscale approach capable of accurately representing the local decaying field. For example, combining deep neural operator (DeepONet) [52] for local surrogate modeling with global FEM solver has recently shown a promising solution to increase the local simulation accuracy [53]. However, it is worth noting that our domain integration method effectively captures the long-range effect of the hole on the loading curve (Fig.2(d)). Our tests of different material parameters, hyperelastic models, and sample geometries show considerable accuracy in the stress contour away from the inhomogeneities. Therefore, we believe our PINNs framework demonstrates noteworthy precision in assessing both the displacement field and the loading history, and the localized stress concentration will have negligible effects on parameter identification in the inverse problem, to be shown in the following subsections.
3.2 Identify constitutive parameters for the prototype problem